# Learning Versatile Humanoid Manipulation with Touch Dreaming

### 14 Apr 2026 v1

@@ V Conclusions
In this work, we study dexterous, contact-rich humanoid manipulation and present an integrated system that combines a robust whole-body controller for stable execution, a real-world VR-based data collection pipeline for humanoid demonstrations, and a touch-aware policy learning framework. We propose Humanoid Transformer with Touch Dreaming (HTD), a multimodal encoder–decoder Transformer that models touch as a core modality and augments single-stage behavioral cloning with future hand-joint-force prediction and future tactile-latent prediction. By supervising future tactile prediction with an EMA target encoder, HTD learns stable and contact-aware latent representations without requiring a separate tactile pretraining stage.
Across five real-world tasks, Insert-T, Book Organization, Towel Folding, Cat Litter Scooping, and Tea Serving, HTD consistently outperforms decoder-only ACT baselines, achieving about 90.9% relative improvement in average success rate over the stronger ACT variant. Our ablations further show that latent tactile dreaming is more effective than raw tactile prediction, with Dream Latent Tactile yielding a 30% relative gain in success rate over Dream Raw Tactile. Together, these results suggest that combining robust whole-body execution, scalable humanoid data collection, and predictive touch-centered learning provides a practical path toward more reliable humanoid manipulation under frequent and complex contact changes.
@@ V-A Lower-Body Controller Details
We provide additional details on the command ranges and domain randomization parameters used in training the lower-body policy.
Command ranges and domain randomization.
Command signals are uniformly sampled within predefined ranges (Table III), where $\bm{v}=[v_{x},v_{y}]^{T}$ denotes the base velocity commands, $\bm{rpy}=[\phi_{\text{torso}},\theta_{\text{torso}},\psi_{\text{torso}}]$ denotes the torso orientation commands, and $h$ is the base height command. To improve sim-to-real transferability, we apply domain randomization across physics parameters (Table IV).
@@
Achievable control ranges.
@@
TABLE V: Maximum stable controllable ranges of our WBC policy in simulation.

### 27 Apr 2026 v2

@@ V Conclusions
In this work, we study dexterous, contact-rich humanoid manipulation and present an integrated system that combines a robust lower-body controller for stable whole-body execution, a real-world VR-based data collection pipeline for humanoid demonstrations, and a touch-aware policy learning framework. We propose Humanoid Transformer with Touch Dreaming (HTD), a multimodal encoder–decoder Transformer that models touch as a core modality and augments single-stage behavioral cloning with future hand-joint-force prediction and future tactile-latent prediction. By supervising future tactile prediction with an EMA target encoder, HTD learns stable, contact-aware latent representations without requiring a separate tactile pretraining stage.
Across five real-world tasks, Insert-T, Book Organization, Towel Folding, Cat Litter Scooping, and Tea Serving, HTD consistently outperforms decoder-only ACT baselines, achieving 90.9% relative improvement in average success rate over the stronger ACT variant. Our ablations further show that latent tactile dreaming is more effective than raw tactile prediction, with Dream Latent Tactile yielding a 30% relative gain in success rate over Dream Raw Tactile. Together, these results suggest that combining robust whole-body execution, scalable humanoid data collection, and predictive touch-centered learning provides a practical path toward more reliable humanoid manipulation under frequent and complex contact changes.
@@ V-A Lower-Body Controller Details
We provide additional details on the command ranges and domain randomization parameters used in training the lower-body controller.
Command ranges and domain randomization
Command signals are uniformly sampled from predefined ranges (Table III), where $\bm{v}=[v_{x},v_{y}]^{T}$ denotes the base velocity commands, $\bm{rpy}=[\phi_{\text{torso}},\theta_{\text{torso}},\psi_{\text{torso}}]$ denotes the torso orientation commands, and $h$ is the base height command. To improve sim-to-real transferability, we apply domain randomization across physics parameters (Table IV).
@@
Achievable control ranges
@@
TABLE V: Maximum stable controllable ranges of our LBC policy in simulation.